Sawaya Co., Ltd., of Japan is a manufacturing company whose total factory overhead costs fluctuate considerably from year to year according to increases and decreases in the number of direct labor-hours worked in the factory. Total factory overhead costs at high and low levels of activity for recent years are given below:

Level of Activity

Low High
  Direct labor-hours 69,600 92,800
  Total factory overhead costs $ 261,360 $ 298,480

The factory overhead costs above consist of indirect materials, rent, and maintenance. The company has analyzed these costs at the 69,600-hour level of activity as follows:

   
  Indirect materials (variable) $ 76,560
  Rent (fixed) 123,000
  Maintenance (mixed) 61,800
  Total factory overhead costs $ 261,360

To have data available for planning, the company wants to break down the maintenance cost into its variable and fixed cost elements.

Required:
1.

Estimate how much of the $298,480 factory overhead cost at the high level of activity consists of maintenance cost. (Hint: To do this, it may be helpful to first determine how much of the $298,480 consists of indirect materials and rent. Think about the behavior of variable and fixed costs!) (Do not round intermediate calculations.)

     

2.

Using the high-low method, estimate a cost formula for maintenance. (Do not round intermediate calculations. Round "Variable cost element" to 2 decimal places.)

     
       

3.

What total factory overhead costs would you expect the company to incur at an operating level of 76,560 direct labor-hours? (Do not round intermediate calculations.)
Answer #1

If you like my explanation, please give a "thumbs up"

Final answers are marked in yellow.

A B C D E F G 1 2 Direct labor-hours Total factory overhead costs 69600 261360 92800 298480 3 st 5 Hour 69600 6 7 76560 8 123

A B C D E F 26 27 Variable cost element Fixed cost element 0.5 27000 28 29 30 Y= 27000 + 0.5 X 31 32 33 3 Activity 76560 34 3

For calculation ref:

A B С D E TI 1 Direct labor-hours Total factory overhead costs 69600 261360 92800 298480 5 Hour 69600 2 3 4 5 6 7 8 9 Indirec

A B с D E F 26 27 Variable cost element Fixed cost element =D25/C25 =D24-C24*C27 28 29 30 Y= 1=C28 + =C27 X 76560 31 32 33 3